Number 142 is divided into two parts such that when the greater part is divided by the smaller, the quotient is 3 and the remainder is 14. What are these parts?
Answer by jjthomson83(4)

Let x and y be the greater and smaller numbers respectively
x+y=142
3y+14=x ie quotient × divisor +remainder= dividend
divisor=y, dividend=x , remainder=14 and quotient=3
x+y=142
x-3y=14
4y=128
y=32
x=142-32
=110
hence x=110 and y=32